OpenStudy (anonymous):

find the slope of the curve when theta= pi/2

OpenStudy (anonymous):

from 0 to pi

hartnn (hartnn):

u know derivatives?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2 cos (2x)

hartnn (hartnn):

finding slope means finding derivative at a point

hartnn (hartnn):

so find dr/dtheta and then put theta =pi/2

OpenStudy (anonymous):

would it be -2?

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

and thats it?

hartnn (hartnn):

yup -2 is the slope

OpenStudy (anonymous):

can someone confirm that?

OpenStudy (anonymous):

looks fine to me

OpenStudy (anonymous):

okay thanks.

OpenStudy (anonymous):

if you graph the polar curve \(\large \theta=\frac{\pi}{2} \), don't you get a vertical line? and what's the slope of a vertical line?

OpenStudy (anonymous):

thats the curve

OpenStudy (anonymous):

oh, ok...